OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = one halfx + 3. y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7

OpenStudy (johnweldon1993):

perpendicular again remember is the negative reciprocal of the slope..so that of 1/2...is -2/1 or just -2 is your slope.. now use the point slope formula again y - (-1) = -2(x - (-3)) y + 1 = -2(x+3) y + 1 = -2x - 6 y = -2x - 7 check -1 = -2(-3) - 7 -1 = 6 - 7 -1 = -1? Yes it does so it works :)